A body starts from rest and moves with a uniform acceleration of 2 m/s 2 in a

straight line.

a) What is the velocity after 5 s

b) How far has it travelled in this time?

c) When will it be 100m from its starting point?

A car accelerates from 4 m/s to 20 m/s in 8 s. How far does it travel in this time?

Differentiate between scalar and vector quantities.

Two forces of 5N and 12N act at right angles at the same point on an object. Find

by scale drawing and calculation the resultant force which is equal in magnitude

and direction to the combined effect of the two forces.

Using a scale of 1 cm to represent 10 N find the size and direction of the resultant

of forces of 30 N and 40 N acting at: a) right angles to each other, b) 60 0 to each

other.


1
Expert's answer
2022-04-14T15:46:24-0400

1.

a) "v=at=10~\frac ms,"

b) "s=\frac{at^2}2=25~m,"

c)

"t'=\sqrt\frac{2s'}a=10~s."

2.

"s=\frac{(v_2-v_1)t}2=64~m."

3.

A scalar quantity has only magnitude, but no direction. Vector quantity has both magnitude and direction. 

4.

"F=\sqrt{F_1^2+F_2^2}=13~N."

5.

a) "F=\sqrt{F_1^2+F_2^2}=50~N,"

b) "F=\sqrt{F_1^2+F_2^2-2F_1F_2\cos 60°}=36~N."